2-2×2+2=?? Please help me
Answer by ikleyn(52781)

.
2-2×2+2=?? Please help me
~~~~~~~~~~~~~~~~~~~~~~~~~~~

2 - 2×2 + 2 = 2 - 4 + 2 = 0.